# How Focused Are LLMs? A Quantitative Study via Repetitive Deterministic Prediction Tasks

### v1

@@ 1 Introduction
However, this paradigm does not directly extend to large language models (LLMs). Many LLM applications, such as machine translation, summarization, or dialogue, are inherently sequence-to-sequence generation problems. In these settings, the model generates text probabilistically, token by token, sampling from a distribution of plausible continuations. Crucially, there is rarely a single “correct” output: multiple valid translations, summaries, or conversational responses may exist. Consequently, simple exact-match accuracy is insufficient for evaluation, and researchers rely instead on a suite of task-specific metrics (e.g., BLEU, ROUGE, perplexity, or human judgment) (Papineni et al., 2002; Lin, 2004; Liang et al., 2022) that capture fluency, coherence, and relevance.
Yet this limitation is context-dependent. In scientific domains—including mathematics, physics, and computer science—the nature of the task often changes. Here, many problems admit a unique correct solution: a definite numerical value, a closed-form expression, or a precise code output. For instance, the question “what is 2342342+34423422342342+3442342” has only one acceptable answer, and the model’s response must exactly match the correct sequence of digits to be considered correct (Cobbe et al., 2021; Hendrycks et al., 2021; Lewkowycz et al., 2022). In such applications, the probabilistic flexibility of LLMs does not diminish the importance of exact correctness. The evaluation once again reduces to a binary criterion: either the model outputs the correct sequence or it does not.
@@
LLMs have achieved remarkable progress across a wide range of tasks, yet their reliability on structured, multi-step reasoning remains poorly understood (Wei et al., 2022; Havrilla et al., 2024). While in generative modeling the challenge has often been framed as a tradeoff between diversity and realism (Astolfi et al., 2024), multi-step reasoning tasks demand consistent correctness across sequential subtasks. Success in these settings is inherently factorizable: the probability of solving the overall problem is the product of per-step accuracies, making performance highly sensitive to individual step error rates. This exposes whether models can exploit compositional structure rather than treating complex problems monolithically (Lake and Baroni, 2018).
To probe this challenge, we introduce a minimal and fully deterministic evaluation class where each problem instance admits a unique ground-truth solution and complexity can be tuned by a single parameter $N$. We present three complementary benchmarks: (i) cyclic letter replacement task, a purely symbolic problem over a finite alphabet that isolates stepwise error propagation without memory; (ii) integer addition, which tests the handling of local carry rules and long-range carry propagation across digits; and (iii) Pauli string multiplication, which requires tracking both local symbol rules and a global phase factor, thereby testing the model’s ability to maintain a lightweight memory across many steps All three benchmarks share the key property that overall complexity decays multiplicatively with sequence length $N$, providing a clean probe of cumulative error accumulation in deterministic reasoning (Wang et al., 2022; Liu et al., 2023; Turpin et al., 2023; Wan et al., 2025a).
Figure 1: Illustration of the experimental setup and theoretical framework. The left panel shows examples of deterministic sequence prediction tasks (e.g., arithmetic problems) used to evaluate large language models (LLMs), where each token’s correctness is represented by a binary variable (Ising spin). Token errors are modeled as Ising spins interacting through all-to-all random couplings, capturing how correlations and noise propagate during sequence generation. The right panel displays the typical behavior of the Sequence Accuracy Rate (SAR) as a function of sequence length $N$: SAR remains high for short sequences, then drops sharply beyond a characteristic crossover scale $N_{*}$, forming the accuracy cliff, which is nicely reproduced by the statistical-mechanical model.
@@ 3.1 Empirical Formula
where $\beta>0$ denotes a constant per-token error rate. However, as we will show in the following, the empirical results deviate significantly from this simple picture. Instead of a smooth exponential decay, SAR exhibits a distinct crossover behavior: it remains nearly perfect for short sequences and then undergoes a rapid drop to zero around a characteristic length scale $N_{*}$.
@@
where $\beta_{0}$ quantifies the intrinsic error rate, and $\alpha$ represents the error accumulation factor, which characterizes how local errors compound as the sequence length increases. When $\alpha=1$, the model reduces to the standard exponential decay expected from independent token errors. For $\alpha>1$, however, the effective error rate $\beta(N)$ grows exponentially with $N$ even when $\beta_{0}$ is small, leading to a rapid, catastrophic drop in accuracy once $\beta(N)$ becomes of order one. The onset of this accuracy cliff occurs at a characteristic sequence length
$N_{*}\sim\frac{\log(1/\beta_{0})}{\log\alpha},$ (21)
@@ 3.2 Hypothesis and Statistical Model
where $h$ denotes an external field reflecting the baseline bias from the prompt toward correct generation, and $J_{ij}$ represent effective couplings between token correctness variables. In the independent-token limit, $J_{ij}=0$ and the model will predict the exponential decay SAR following Eq. (19). However, experimental results described by Eq. (20) clearly indicate deviations from this independent picture, implying the presence of significant correlations among token outcomes.
@@ 4 Related Works
Single-pass reasoning and capacity limits. Recent research has shown that even when models can locally apply correct reasoning steps, global accuracy collapses as reasoning chains lengthen. This phenomenon parallels information-theoretic capacity limits of single forward passes. The Fano-style upper bound framework  (Wan et al., 2025a) formalizes this collapse for multi-hop question answering (MHQA), proving that once task complexity exceeds the model’s per-pass information capacity, accuracy must drop sharply—mirroring our empirical SAR crossover. Their work interprets the reasoning chain as a noisy communication channel, bounded by mutual information, while our SAR analysis arrives at the same breakdown curve from a statistical-physics viewpoint, emphasizing the all-to-all error propagation in the attention architecture as the key underlying mechanism. Together they highlight the inevitability of reliability loss in single-pass reasoning under finite model capacity.
Prompting paradigms and decomposition. A wide range of prompting methods—Direct prompting, Chain-of-Thought (CoT) (Wei et al., 2022; Kojima et al., 2022), ReAct (Yao et al., 2023), Plan-and-Solve (Wang et al., 2023), and Self-Ask (Press et al., 2023)—operate in the single-pass regime, assuming the full reasoning chain fits within one generation. By contrast, multi-call frameworks such as Self-Refine (Madaan et al., 2023), InfoQA (Wan et al., 2025a), and agentic systems for programming or reasoning (Li et al., 2024; Qian et al., 2024; Kim et al., 2024) mitigate this bottleneck by decomposing reasoning into smaller, verifiable subcalls. Our SAR-based analysis provides a quantitative explanation of why such multi-call methods succeed: by resetting the error accumulation process, they effectively bound per-step complexity below the model’s capacity threshold, preventing the exponential or super-exponential degradation seen in single-pass reasoning.
Self-checking, verification, and robust reasoning. Other orthogonal strategies—such as self-consistency sampling, debate-style verification, and tool-augmented reasoning—attempt to mitigate compounding errors by introducing redundancy or external validation (Yao et al., 2023; Xie et al., 2025; Wan et al., 2025b). SAR offers a unified diagnostic lens for such methods: improved reliability corresponds to reduced intrinsic error rate $\beta_{0}$, while better verification or decomposition corresponds to smaller accumulation factor $\alpha$. Thus, SAR quantitatively links algorithmic interventions to empirical stability regimes.
@@ Appendix A Success Rate Analysis of Pauli String test
Because this factorizable Pauli multiplication problem constitutes the simplest possible toy model of a multi-step workflow, it serves as a probe into how LLMs manage reasoning, error accumulation, and memory in more complex agentic tasks.
@@ B.3 Replica trick and disorder average
From Eq.˜43, it is straightforward to show that

### v2

@@ 1 Introduction
However, this paradigm does not directly extend to large language models (LLMs). Many LLM applications, such as machine translation, summarization, or dialogue, are inherently sequence-to-sequence generation problems. In these settings, the model generates text probabilistically, token by token, sampling from a distribution of plausible continuations. Crucially, there is rarely a single “correct” output: multiple valid translations, summaries, or conversational responses may exist. Consequently, simple exact-match accuracy is insufficient for evaluation, and researchers rely instead on a suite of task-specific metrics (e.g., BLEU, ROUGE, perplexity, or human judgment) (papineni-etal-2002-bleu; lin-2004-rouge; liang2022holistic) that capture fluency, coherence, and relevance.
Yet this limitation is context-dependent. In scientific domains—including mathematics, physics, and computer science—the nature of the task often changes. Here, many problems admit a unique correct solution: a definite numerical value, a closed-form expression, or a precise code output. For instance, the question “what is 2342342+34423422342342+3442342” has only one acceptable answer, and the model’s response must exactly match the correct sequence of digits to be considered correct (cobbe2021training; hendrycks2021measuring; lewkowycz2022solving). In such applications, the probabilistic flexibility of LLMs does not diminish the importance of exact correctness. The evaluation once again reduces to a binary criterion: either the model outputs the correct sequence or it does not.
@@
LLMs have achieved remarkable progress across a wide range of tasks, yet their reliability on structured, multi-step reasoning remains poorly understood (wei2022chain; havrilla2024glore). While in generative modeling the challenge has often been framed as a tradeoff between diversity and realism (astolfi2024consistency), multi-step reasoning tasks demand consistent correctness across sequential subtasks. Success in these settings is inherently factorizable: the probability of solving the overall problem is the product of per-step accuracies, making performance highly sensitive to individual step error rates. This exposes whether models can exploit compositional structure rather than treating complex problems monolithically (lake2018generalization).
To probe this challenge, we introduce a minimal and fully deterministic evaluation class where each problem instance admits a unique ground-truth solution and complexity can be tuned by a single parameter $N$. We present three complementary benchmarks: (i) cyclic letter replacement task, a purely symbolic problem over a finite alphabet that isolates stepwise error propagation without memory; (ii) integer addition, which tests the handling of local carry rules and long-range carry propagation across digits; and (iii) Pauli string multiplication, which requires tracking both local symbol rules and a global phase factor, thereby testing the model’s ability to maintain a lightweight memory across many steps All three benchmarks share the key property that overall complexity decays multiplicatively with sequence length $N$, providing a clean probe of cumulative error accumulation in deterministic reasoning (wang2022self; liu2023lost; turpin2023language; wan2025fano).
Figure 1: Illustration of the experimental setup and theoretical framework. The left panel shows examples of deterministic sequence prediction tasks (e.g., arithmetic problems) used to evaluate large language models (LLMs), where each token’s correctness is represented by a binary variable (Ising spin). Token errors are modeled as Ising spins interacting through all-to-all random couplings, capturing how correlations and noise propagate during sequence generation. The right panel displays the typical behavior of the Sequence Accuracy Rate (SAR) (i.e. the probability for all tokens to be correct in a output sequence) as a function of the sequence length $N$: SAR remains high for short sequences, then drops sharply beyond a characteristic crossover scale $N_{*}$, forming the accuracy cliff, which is nicely reproduced by a spin-glass statistical-mechanical model.
@@ 3.1 Empirical Formula
where $\beta>0$ denotes a constant per-token error rate. However, as we have seen in the above experiments, the empirical results deviate significantly from this simple picture. Instead of a smooth exponential decay, SAR exhibits a distinct crossover behavior: it remains nearly perfect for short sequences and then undergoes a rapid drop to zero around a characteristic length scale $N_{*}$.
@@
where $\beta_{0}>0$ quantifies the intrinsic error rate, and $\alpha$ represents the error accumulation factor, which characterizes how local errors compound as the sequence length increases. When $\alpha=1$, the model reduces to the standard exponential decay expected from independent token errors. For $\alpha>1$, however, the effective error rate $\beta(N)$ grows exponentially with $N$ even when $\beta_{0}$ is small, leading to a rapid, catastrophic drop in accuracy once $\beta(N)$ becomes of order one. The onset of this accuracy cliff occurs at a characteristic sequence length
$N_{*}\simeq 1+\frac{\log(1/\beta_{0})}{\log\alpha},$ (21)
@@ 3.2 Hypothesis and Statistical Model
where $h$ denotes an external field reflecting the baseline bias from the prompt toward correct generation, and $J_{ij}$ represent effective couplings between token correctness variables. In the independent-token limit, $J_{ij}=0$ and the model will predict the exponential decay SAR following Eq. (19). However, experimental results described by Eq. (20) clearly indicate deviations from this independent picture, implying the presence of correlations among token outcomes.
@@
4 Implication: Divide-and-Conquer Strategy
The empirical scaling law of Eq. (20) not only provides a qualitative description of how the accuracy inevitably degrades with the output sequence length $N$ for LLMs, but also offers a constructive implication: when $\alpha>1$, the reliability of long-sequence generation can be improved by adopting a divide-and-conquer strategy.
In this context, divide-and-conquer refers to partitioning a sequence-generation task of total length $N$ into $k$ sub-tasks of length $N/k$, prompting the model separately for each sub-task, and then merging the partial outputs to form the complete sequence. The intuition is that the observed accuracy cliff originates from correlated error propagation among tokens: when $\alpha>1$, the effective per-token error rate grows exponentially with $N$, driven by all-to-all dependencies in the attention mechanism. Dividing the sequence into shorter fragments effectively cuts the correlation loops, preventing the catastrophic error accumulation.
We introduce a multiplicative overhead factor $\theta_{N,k}\in(0,1]$ to account for the reduction of SAR due to the additional operations involved in applying the divide-and-conquer strategy, such as segmentation, formatting, boundary handling, and content copying. $\theta_{N,k}$ is expected to only depend on $N$ and $k$ weakly. The overall success probability under divide-and-conquer is then
$\mathrm{SAR}_{\mathrm{DC}}(N,k)=\theta_{N,k}\,\big[\mathrm{SAR}(N/k)\big]^{k}=\theta_{N,k}\,\exp\!\Big[-\beta_{0}\,N\,\alpha^{\,\frac{N}{k}-1}\Big].$ (28)
Define the logarithmic gain of divide-and-conquer relative to the single-shot strategy as
$\Delta(N,k):=\log\frac{\mathrm{SAR}_{\mathrm{DC}}(N,k)}{\mathrm{SAR}(N)}=\log\theta_{N,k}+\beta_{0}N\Big(\alpha^{\,N-1}-\alpha^{\,\frac{N}{k}-1}\Big).$ (29)
This formulation shows that for $\alpha>1$, the gain term increases rapidly with $N$, indicating that the larger $\alpha$ is, the more advantageous segmentation becomes.
Theorem 1 (Advantage of Divide-and-Conquer).
Let the intrinsic error rate $\beta_{0}>0$ and the correlation amplification factor $\alpha>1$, and fix the integer $k\geq 2$. Under the empirical scaling law $\mathrm{SAR}(N)=\exp[-\beta_{0}N\alpha^{\,N-1}]$, for the segmentation into $k$ equal sub-tasks to yield a positive gain $\Delta(N,k)>0$, it would be sufficient if the sequence length exceeds the following bound
$N\geq N_{\mathrm{DC}}\;=\;1+\frac{1}{\log\alpha}\left(\log\left(1-\frac{2\log\theta_{N,k}}{\beta_{0}}\right)+\frac{\log 2}{1-1/k}\right),$ (30)
meaning that for all $N\geq N_{\mathrm{DC}}$, one has $\mathrm{SAR}_{\mathrm{DC}}(N,k)>\mathrm{SAR}(N)$.
(See App. C for a proof.)
This result provides an explicit criterion for when the divide-and-conquer approach is beneficial. Under this strategy, as evident from Eq. (28), the intrinsic error rate $\beta_{0}$ remains unchanged, while the correlation amplification factor $\alpha$ is effectively rescaled to $\alpha_{\text{eff}}\simeq\alpha^{1/k}$ in the large-$N$ regime. As a consequence, the characteristic accuracy-cliff scale is extended to
$N_{*}\simeq 1+\frac{\log(1/\beta_{0})}{\log\alpha_{\text{eff}}}=1+k\,\frac{\log(1/\beta_{0})}{\log\alpha},$ (31)
growing approximately linearly with $k$, according to Eq. (21). The empirical effectiveness of the divide-and-conquer strategy is demonstrated in Fig.˜8, where the linear-$k$ scaling of accuracy cliff is indeed observed.
In summary, the empirical scaling law does more than describe the accuracy cliff—it also predicts how to mitigate it: segmentation effectively reduces the compounding of correlated errors and extends the reliable length scale of LLMs by roughly a factor of $k$, transforming an observed limitation into a quantitative improvement strategy.
Figure 8: Divide-and-conquer evaluation on the Pauli string multiplication task under both phase-restricted and phase-relaxed criteria, using gemini-2.5-flash and gemini-2.5-pro. The gray k=1k{=}1 curves show the fitted $\mathrm{SAR}(N)$ from Eq.˜20; the extracted $(\alpha,\beta_{0})$ values are then used to plot the theoretical $\mathrm{SAR}(N/k)^{k}$ curves for k=2k{=}2 and k=3k{=}3 (red and blue), assuming no overhead ($\theta_{N,k}=1$). In the phase-restricted case, the pro model closely follows the predicted improvement, while the flash model performs below the theoretical curves, due to additional overhead in divide and combine. In the phase-relaxed case, both models closely follow the predicted improvement, and additionally expand the perfect-accuracy plateau by approximately a factor of $k$ (indicated by the colored dashed lines).
@@ 5 Related Works
Single-pass reasoning and capacity limits. Recent research has shown that even when models can locally apply correct reasoning steps, global accuracy collapses as reasoning chains lengthen. This phenomenon parallels information-theoretic capacity limits of single forward passes. The Fano-style upper bound framework  (wan2025fano) formalizes this collapse for multi-hop question answering (MHQA), proving that once task complexity exceeds the model’s per-pass information capacity, accuracy must drop sharply—mirroring our empirical SAR crossover. Their work interprets the reasoning chain as a noisy communication channel, bounded by mutual information, while our SAR analysis arrives at the same breakdown curve from a statistical-physics viewpoint, emphasizing the all-to-all error propagation in the attention architecture as the key underlying mechanism. Together they highlight the inevitability of reliability loss in single-pass reasoning under finite model capacity.
Prompting paradigms and decomposition. A wide range of prompting methods—Direct prompting, Chain-of-Thought (CoT) (wei2022chain; kojima2022large), ReAct (yao2023react), Plan-and-Solve (wang2023plan), and Self-Ask (press2023measuring)—operate in the single-pass regime, assuming the full reasoning chain fits within one generation. By contrast, multi-call frameworks such as Self-Refine (madaan2023self), InfoQA (wan2025fano), and agentic systems for programming or reasoning (li2024survey; qian2024chatdev; kim2024llm) mitigate this bottleneck by decomposing reasoning into smaller, verifiable subcalls. Our SAR-based analysis provides a quantitative explanation of why such multi-call methods succeed: by resetting the error accumulation process, they effectively bound per-step complexity below the model’s capacity threshold, preventing the exponential or super-exponential degradation seen in single-pass reasoning.
Self-checking, verification, and robust reasoning. Other orthogonal strategies—such as self-consistency sampling, debate-style verification, and tool-augmented reasoning—attempt to mitigate compounding errors by introducing redundancy or external validation (yao2023react; xie2025fire; wan2025cognitive; mccoy2023embers). SAR offers a unified diagnostic lens for such methods: improved reliability corresponds to reduced intrinsic error rate $\beta_{0}$, while better verification or decomposition corresponds to smaller accumulation factor $\alpha$. Thus, SAR quantitatively links algorithmic interventions to empirical stability regimes.
@@ B.3 Replica trick and disorder average
From Eq.˜47, it is straightforward to show that
@@ B.3.1 Small-$J_{0}$ Expansion
Appendix C Proof of Theorem 1
Proof.
Start with the condition
$N\geq N_{\mathrm{DC}}\;=\;1+\frac{1}{\log\alpha}\left(\log\left(1-\frac{2\log\theta_{N,k}}{\beta_{0}}\right)+\frac{\log 2}{1-1/k}\right),$ (67)
with $\alpha>1$, $\beta_{0}>0$, $0<\theta_{N,k}\leq 1$, $k\geq 2$. Each term in $N_{\mathrm{DC}}$ is non-negative,
$\log\alpha>0,\quad\log\left(1-\frac{2\log\theta_{N,k}}{\beta_{0}}\right)\geq 0,\quad\frac{\log 2}{1-1/k}>0,$ (68)
therefore Eq. (67) implies the following weaker conditions:
$N>1,$ (69)
$N>1+\frac{1}{\log\alpha}\log\left(1-\frac{2\log\theta_{N,k}}{\beta_{0}}\right)>1+\frac{1}{\log\alpha}\log\left(-\frac{2\log\theta_{N,k}}{\beta_{0}}\right),$ (70)
$N>\frac{1}{\log\alpha}\frac{\log 2}{1-1/k},$ (71)
which are nevertheless more directly useful.
Eq. (71) implies
$\begin{split}&N(1-1/k)\log\alpha>\log 2\\ \Rightarrow\;&\alpha^{N(1-1/k)}>2\\ \Rightarrow\;&\alpha^{-N(1-1/k)}<1/2\\ \Rightarrow\;&(1-\alpha^{N/k-N})>1/2.\end{split}$ (72)
Eq. (70) implies
$(N-1)\log\alpha>\log\left(-\frac{2\log\theta_{N,k}}{\beta_{0}}\right)\Rightarrow\;\frac{1}{2}\beta_{0}\alpha^{N-1}>-\log\theta_{N,k}.$ (73)
Putting Eqs. (69), (72), (73) together, we have
$\begin{split}\beta_{0}N(\alpha^{N-1}-\alpha^{N/k-1})&=\beta_{0}\alpha^{N-1}N(1-\alpha^{N/k-N})\\ &>\beta_{0}\alpha^{N-1}\times 1\times\frac{1}{2}\\ &>-\log\theta_{N,k},\end{split}$ (74)
therefore the logarithmic gain $\Delta(N,k)=\log\theta_{N,k}+\beta_{0}N(\alpha^{N-1}-\alpha^{N/k-1})>0$ is positive. ∎